# TimeMixer++: A General Time Series Pattern Machine for Universal Predictive Analysis

### v1

@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2023), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024), TIDE(Das et al., 2023a), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023), and SCINet(Liu et al., 2022a).
@@
As illustrated in Figure 13 for imputation in Weather and long-term forecasting in ETTm1, TimeMixer++ demonstrates a balance between low memory footprint and training time while achieving competitive MSE scores. In the Weather imputation task, TimeMixer++ has a smaller memory footprint (2.9GB) and moderate training time (395ms) compared to other models like FEDformer, which has higher memory usage (9.8GB) and longer training time (812ms). For long-term forecasting, TimeMixer++ also shows efficiency with the lowest memory footprint (2.1GB) and competitive training time (274ms), while maintaining a low MSE. To be further detailed, TimeMixer++ demonstrates efficiency by transforming 1D time series data into 2D images. This transformation allows the use of parameter-efficient inception block 2D convolutions, which not only enhance computational efficiency but also effectively capture long-range dependencies that are difficult to detect in 1D sequences. This approach results in a balance between low memory footprint and training time while achieving competitive MSE scores, making TimeMixer++ highly suitable for tasks requiring fast and resource-efficient processing.
@@ Appendix F Hyperparamter Sensitivity
Considering overall performance and efficiency, the marginal benefits of increasing $M$ significantly diminish, so we set $M$ to 3. (2) Adding more layers typically reduces MSE, particularly between 1 and 2 layers where the change is most pronounced. For shorter prediction lengths (e.g., predict-96), increasing the number of layers results in more significant performance gains. Considering all factors, we set $L$ to 2. (3) Increasing the selection of Top-K generally leads to a reduction in MSE. For longer prediction lengths (e.g., predict-720), the choice of Top-K has a more substantial impact on the results. Based on the above considerations, we choose $K$ to be 3. (4) As the input length increases, the MSE gradually decreases. Longer input lengths help improve prediction accuracy across all prediction lengths, especially noticeable in the transition from 96 to 192 input length. This indicates that using longer inputs can achieve better predictive performance for TimeMixer++.
@@ Appendix G Error bars
Model TimeMixer++ iTransformer (2023) Confidence
@@
Model TimeMixer++ TimeMixer (2024) Confidence
@@
Model TimeMixer++ TimesMixer (2024) Confidence
@@ Appendix H Full Results
Due to the space limitation of the main text, we place the full results of all experiments in the following: long-term forecasting in Table 16, univariate short-term forecasting in Table 17, multivariate short-term forecasting in Table 18, classification in Table 19 and anomaly detection in Table 20.
Table 16: Unified hyperparameter results for the long-term forecasting task. We compare extensive competitive models under different prediction lengths. Avg is averaged from all four prediction lengths, that $\{96,192,336,720\}$.

### v2

@@ Appendix E Efficiency Analysis
We comprehensively compare the forecasting and imputation in performance, training speed, and memory footprint of the following models: TimeMixer++, iTransformer(Liu et al., 2023a), PatchTST(Nie et al., 2023), TimeMixer(Wang et al., 2024b), TIDE(Das et al., 2023a), Fedformer(Zhou et al., 2022b), TimesNet(Wu et al., 2023), MICN(Wang et al., 2023a), and SCINet(Liu et al., 2022a).
@@
As shown in Figure 14, TimeMixer++ achieves a comparable balance among memory footprint, training time, and performance in both the Weather imputation and long-term forecasting tasks on ETTm1, delivering the MSE scores.
@@ Appendix F Hyperparamter Sensitivity
Considering overall performance and efficiency, the marginal benefits of increasing $M$ significantly diminish, so we can set $M$ from 1 to 3. (2) Adding more layers typically reduces MSE, particularly between 1 and 2 layers where the change is most pronounced. For shorter prediction lengths (e.g., predict-96), increasing the number of layers results in more significant performance gains. (3) Increasing the selection of Top-K generally leads to a reduction in MSE. For longer prediction lengths, the choice of Top-K has a more substantial impact on the results. (4) As the input length increases, the MSE gradually decreases. Longer input lengths help improve prediction accuracy across all prediction lengths, which indicates that using longer inputs may achieve better predictive performance for the TimeMixer++.
@@ Appendix G Error bars
Model TimeMixer++ iTransformer (2023a) Confidence
@@
Model TimeMixer++ TimeMixer (2024b) Confidence
@@
Model TimeMixer++ TimesMixer (2024b) Confidence
@@ Appendix H Full Results
Due to the space limitation of the main text, we place the full results of all experiments in the following: long-term forecasting in Table 16, univariate short-term forecasting in Table 17, multivariate short-term forecasting in Table 18, imputation in Table 19, few-shot learning in Table 20, classification in Table 21 and anomaly detection in Table 22.
Table 16: Full results for the long-term forecasting task. We compare extensive competitive models under different prediction lengths. Avg is averaged from all four prediction lengths, that $\{96,192,336,720\}$.
@@
Table 19: Full results of imputation task across six datasets. To evaluate our model performance, we randomly mask $\{12.5\%,25\%,37.5\%,50\%\}$ of the time points in time series of length 1024. The final results are averaged across these 4 different masking ratios.
Models
TimeMixer++
TimeMixer
iTransformer
PatchTST
Crossformer
FEDformer
TIDE
DLinear
TimesNet
MICN
Autoformer
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.041
0.127
0.072
0.178
0.075
0.177
0.097
0.194
0.0810
0.196
0.048
0.152
0.090
0.210
0.076
0.191
0.049
0.147
0.059
0.170
0.062
0.169
ETTm2
0.024
0.135
0.061
0.166
0.055
0.169
0.080
0.183
0.111
0.219
0.087
0.198
0.169
0.263
0.088
0.198
0.035
0.124
0.109
0.221
0.077
0.181
ETTh1
0.091
0.198
0.152
0.242
0.130
0.213
0.178
0.231
0.201
0.309
0.099 0.225
0.289
0.395
0.174
0.288
0.142
0.258
0.127
0.254
0.148
0.267
ETTh2
0.065
0.157
0.101
0.294
0.125
0.259
0.124
0.293
0.206
0.308
0.262
0.344
0.709
0.596
0.120
0.238
0.088 0.198
0.179
0.290
0.129
0.242
ECL
0.109
0.197
0.142
0.261
0.140
0.223
0.129
0.198 0.125
0.204
0.181
0.314
0.182
0.202
0.080
0.200
0.135
0.255
0.138
0.246
0.141
0.234
Weather
0.049
0.078
0.091
0.114
0.095
0.102
0.082
0.149
0.150
0.111
0.064
0.139
0.063
0.131
0.071
0.107
0.061 0.098
0.075
0.126
0.066
0.107
Models
TimeMixer++
TimeMixer
iTransformer
TiDE
Crossformer
DLinear
PatchTST
TimesNet
FEDformer
Autoformer
Stationary
ETSformer
LightTS
Informer
Reformer
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
0.517
0.512
0.613
0.52
0.51
0.597
0.589
0.535
0.645
0.558
0.691
0.600
0.633
0.542
0.869
0.628
0.639
0.561
0.702
0.596
0.915
0.639
1.180
0.834
1.375
0.877
1.199
0.809
1.249
0.833
ETTh2
0.379
0.391
0.402
0.433
0.455
0.461
0.395 0.412
0.428
0.447
0.605
0.538
0.415
0.431
0.479
0.465
0.466
0.475
0.488
0.499
0.462
0.455
0.894
0.713
2.655
1.160
3.872
1.513
3.485
1.486
ETTm1
0.398
0.431
0.487
0.461
0.491
0.516
0.425
0.458
0.462
0.489
0.411
0.429
0.501
0.466
0.677
0.537
0.722
0.605
0.802
0.628
0.797
0.578
0.980
0.714
0.971
0.705
1.192
0.821
1.426
0.856
ETTm2
0.291
0.351
0.311
0.367
0.375
0.412
0.317
0.371
0.343
0.389
0.316
0.368
0.296
0.343
0.320
0.353
0.463
0.488
1.342
0.930
0.332
0.366
0.447
0.487
0.987
0.756
3.370
1.440
3.978
1.587
Weather
0.241
0.271
0.242
0.281
0.291
0.331
0.249
0.291
0.267
0.306
0.241
0.283
0.242
0.279
0.279
0.301
0.284
0.324
0.300
0.342
0.318
0.323
0.318
0.360
0.289
0.322
0.597
0.495
0.546
0.469
ECL
0.168
0.271
0.187
0.277
0.241
0.337
0.196
0.289
0.214
0.308
0.180
0.280
0.180 0.273
0.323
0.392
0.346
0.427
0.431
0.478
0.444
0.480
0.660
0.617
0.441
0.489
1.195
0.891
0.965
0.768